# DeSocial: Blockchain-based Decentralized Social Networks

### v1

@@ 1 Introduction
Social network learning algorithms have become a central tool for modeling and predicting user behavior in social networks JODIE ; MetaDyGcn ; TGAT ; SGFormer ; ROLAND , enhancing content search, advertising and recommendation, and improve user experience across multiple platforms SML ; IGC ; xia2023automated ; MetaDyGcn ; SGC ; LightGCN ; wang2021inductive . Despite impressive advances in graph-based recommendations wu2020comprehensive ; huang2023position ; xu2024rethinking ; yang2023dgrec ; yu2023towards ; xu2023neural ; TGOnline ; wang2022rete , the main applications are still deeply rooted in the Web 2.0 paradigm of centralized platforms with exclusive control over user data and model deployment. The platform itself selects the predictive algorithms, trains the models, and provides recommendations to users without transparency or user involvement user_control1 ; ge2024survey . As a result, users are forced to accept predictions derived from a fixed model pipeline, regardless of whether the prediction fits their personal preferences or the structural context. This model limits personalization and does not reflect the diversity of the user’s local environments, especially in large graphs. Moreover, relying on a single model limits expressive power, even if that model is state-of-the-art. These challenges point to the need for frameworks that allow users to play a more active role in the algorithm selection and prediction process, especially in decentralized environments that effectively support model diversity and personalization.
@@
By contrast, blockchain as a Web 3.0 technology is a promising alternative jeong2025navigating ; nakamoto2008bitcoin ; wood2014ethereum ; yakovenko2018solana ; binance_us . With their emphasis on decentralization mlika2024blockchain ; jeong2025navigating , transparency zou2019smart ; li2023demystifying and verifiable interactions qiu2024lido ; liu2019survey , blockchain-based systems offer a new paradigm for social, transactional and economic behavior that is inherently more user-centric and transparent wan2024web3 ; wohlhieter2021bitclout ; liu2023tokenomic ; myshell_docs . Despite their rapid development, blockchain systems are not yet fully integrated with modern artificial intelligence approaches, especially in social network applications, where such synergies could be most impactful jeong2025navigating ; wan2024web3 . Therefore, there is a significant gap in applying graph learning algorithms to decentralized social prediction tasks, which provides new capabilities for building more personalized, user-driven, and decentralized social network systems.
To overcome this gap, we propose DeSocial framework that allows users to regain control over the forecasting process on their own behalf. In DeSocial , each user has their rights to select the most favorable prediction model from a library of graph learning networks (e.g., MLP MLP , GCN GCN , GAT GAT , GraphSAGE graphsage , or SGC SGC ). This selection is based on neighborhood sampling evaluation to ensure that the selected model best captures the user’s local environment. To ensure reliability, DeSocial applies a smart contract that is transparent to every nodes: a group of validators are selected to run their models independently, evaluate the link prediction query, and make the decision together by a majority vote. This approach not only increases the robustness of the predictions, but also removes single-point algorithmic control, aligning outcomes with both user intent and system integrity. Our contributions can be summarized as follows:
@@ 2.1 Graph Learning for Social Networks
Existing graph learning methods for social networks are advanced, including graph neural networks (GNN) GCN ; GAT ; graphsage ; SGC ; GIN ; GPS ; LightGCN ; JODIE ; DySAT ; evolvegcn ; wang2021inductive , meta-learning frameworks SML ; IGC ; ROLAND ; MetaDyGcn ; TGOnline , Transformer-based SGFormer ; NodeFormer ; ying2021transformers , and large language model (LLM)-based GraphPrompter ; GraphGPT ; LLaGA ; LLMRec ; InstructGLM ; he2023harnessing ; yu2025graph2text ; PromptGFM , most of which are designed for centralized environments, where the full graph structure and node features are aggregated on a single server and optimized jointly. In contrast, our setting assumes a completely decentralized environment where each user node has full access to the local view of the social network structure, but raw data cannot be shared across nodes zhang2021federated ; qi2022fairvfl ; meng2021cross ; qin2024blockdfl . Furthermore, the storage and computational capabilities of each node are limited, making it infeasible to deploy large-scale models such as LLM-based encoders li2023graphtransformer ; GraphGPT ; 10697304 ; ren2024llmgraphsurvey or even invoke external LLM APIs xu2024slmrec ; liu2025uncovering ; mei2024aios ; xu2025mem ; zhang2024mabc ; zhang2023building ; jin2025massive ; xu2024slmrec ; jin-etal-2025-exploring ; jin2024disentangling . This practical limitation leads us to limit each node to implementing a lightweight graph foundation. Therefore, we selected five representative and well-studied models: MLP MLP , GCN GCN , GAT GAT , GraphSAGE graphsage , and SGC SGC , as candidate backbones for our framework. These backbones achieve a practical balance between representational power and computational efficiency, which is sufficient to test the effectiveness of model selection and consensus in decentralized graph learning.
@@ 2.2 Blockchain Consensus Mechanism
Blockchain consensus mechanisms enable distributed agreement without central control mingxiao2017review ; nakamoto2008bitcoin ; wood2014ethereum . The original Bitcoin blockchain relied on Proof-of-Work (PoW) nakamoto2008bitcoin consensus, where the longest chain determines the accepted state. While PoW provides an open network, it is energy intensive and only provides probabilistic finiteness. More relevant to our context are the consensus protocols Proof-of-Stake (PoS) wood2014ethereum ; yakovenko2018solana ; binance_us , Proof-of-Authority PoA ; binance_us , and Byzantine Fault Tolerance (BFT) aptos2022whitepaper ; mazieres2015stellar ; sui2022whitepaper , which achieve finality by voting among selected validators. However, these methods do not achieve complete decentralization fahim2023blockchain ; jeong2024centralized , or a large number of validators are selected to participate in the verification, which greatly reduces the operation speed of the network yin2019hotstuff . These concerns of scalability, centralization, and running efficiency bring challenges to blockchain-based decentralized social network algorithm design. Therefore, a limited set of validators are selected to participate in the consensus for each verification in DeSocial , which guarantees the efficiency, robustness, and complete decentralization of the network.
@@ 2.3 Ensemble Learning and Majority Voting
Ensemble learning that combines multiple models to improve predictions is well-recognized in machine learning dong2020survey ; breiman1996bagging ; freund1997decision ; mienye2022survey ; roh2024levi . In graph learning, ensemble methods help mitigate overfitting to specific topologies or noisy subgraphs, especially when models are trained on different views of the graph or initialized with different parameters duan2024graph ; zhang2024e2gnn ; chen2022ensemble . Voting is essential in blockchain consensus process because it enables nodes to collectively agree on the validity of transactions xu2021multi . Similarly, voting serves as a mechanism for aggregating decisions from multiple nodes in decentralized graph learning, ensuring that predictions are agreed upon through collective decision-making rather than centralized authority. Common voting strategies includes soft voting wen2024black ; ye2025pilot , where models make probabilistic predictions and use averages, and hard voting franklin2011crowddb ; wen2024black ; ye2025pilot , where each model casts a binary vote and the final outcome is the majority. DeSocial adopts hard voting since it aligns naturally with blockchain consensus mechanisms, where validations are made through majority approval. Voting-based decisions are also more robust and less vulnerable to manipulation.
@@ 5.1 Experimental Setups
We use four real-world temporal graph datasets in the scope of Web 2.0 (UCI poursafaei2022towards , Enron zhang2024dtgb , and GDELT zhang2024dtgb ) and Web 3.0 (Memo-Tx zuo2023set ). In order to evaluate the performance of DeSocial , we used the five most classic centralized models (MLP MLP , GCN GCN , GAT GAT , GraphSAGE graphsage , and SGC SGC ) in the field of social network learning as baselines for comparison. Detailed information for the datasets and baselines is described in Appendix A.1 and Appendix A.2, respectively.
@@ 5.2 Evaluation Metrics
Given the decentralized nature of our problem formulation, we employ evaluation metrics tailored to assess the effectiveness of DeSocial in decentralized environments, enlightened by the validation of the double spending problem on blockchain liu2024two ; kumar2023review . That is, they lack access to the underlying intent of a transaction and can only judge based on structural context. This motivates our use of Acc@K, a set of relative metrics that evaluates whether the predicted positive link ranks higher than its sampled alternatives.
Given a link prediction task that each test case consists of one positive edge and $K-1$ negative edges, Acc@K indicates the probability that the model assigns the highest score to the positive edge among the $K$ candidates. We adopt $K\in\{2,3,5\}$ to evaluate the performance of decentralized graph learning. The greater $K$ is, the harder evaluation tasks will be. We follow  zhang2024dtgb for the randomized negative sampling method. Appendix A.3 provides our experiment setups and implementation details.
@@ 5.3 Impacts of the Personalized Algorithms
Figure 4 reports the performance of five centralized baselines, random selection, simple rule-based selection, and our personalized algorithm selection method DeSocial PA across multiple datasets and evaluation metrics. In hybrid settings, each validator uses a distinct backbone. Random selection represents that every node selects an algorithm in $\mathcal{F}$. Simple rule-based selection represents that every node selects an algorithm in $\mathcal{F}$ based on its two features (degree and clustering) of the local structure.  Table 1 and  Table 2 shows the detailed statistics of the centralized baselines and DeSocial PA. Appendix A.5 shows the simple rule of algorithm selection.
@@ 5.5 Efficiency Analysis
We use Ganache to capture blockchain-side latency and execution cost in decentralized social prediction processes, avoiding the complexity of deploying to a full public ETH testnet. For decentralized methods, runtime is measured from a single user’s perspective. That is, starting from a social link request and ending to the decision aggregated. Appendix B shows the details of calculating the run time of each decentralized method.
As shown in  Table 3, the amortized runtime of DeSocial is not significantly impacted by the integration of the ETH local development chain. Our experiments are conducted in a single machine simulation where both graph training and blockchain interactions are executed serially. Appendix A.3 shows the details of the computation device. In a real ETH network deployment with multiple participating users, the system would better exploit parallelism and further improve the efficiency.
@@ A.3 Implementation Details
Code Availability. We have included the implementation code in our supplementary zip file. Although we originally planned to provide an anonymous Github repository, we decided that the supplementary materials already suffice for ensuring reproducibility.

### v2

@@ 1 Introduction
Social network learning algorithms have become a central tool for modeling and predicting user behavior in social networks [26, 77, 69, 66, 83], enhancing content search, advertising and recommendation, and improve user experience across multiple platforms [91, 6, 68, 77, 64, 14, 59]. Despite impressive advances in graph-based recommendations [67, 16, 75, 78, 84, 72, 57, 58], the main applications are still deeply rooted in the Web 2.0 paradigm of centralized platforms with exclusive control over user data and model deployment. The platform itself selects the predictive algorithms, trains the models, and provides recommendations to users without transparency or user involvement [17, 12]. As a result, users are forced to accept predictions derived from a fixed model pipeline, regardless of whether the prediction fits their personal preferences or the structural context. This model limits personalization and does not reflect the diversity of the user’s local environments, especially in large graphs. Moreover, relying on a single model limits expressive power, even if that model is state-of-the-art. These challenges point to the need for frameworks that allow users to play a more active role in the algorithm selection and prediction process, especially in decentralized environments that effectively support model diversity and personalization.
@@
By contrast, blockchain as a Web 3.0 technology is a promising alternative [19, 42, 63, 76, 2]. With their emphasis on decentralization [40, 19], transparency [93, 29] and verifiable interactions [47, 30], blockchain-based systems offer a new paradigm for social, transactional and economic behavior that is inherently more user-centric and transparent [56, 62, 31, 41]. Despite their rapid development, blockchain systems are not yet fully integrated with modern artificial intelligence approaches, especially in social network applications, where such synergies could be most impactful [19, 56]. Therefore, there is a significant gap in applying graph learning algorithms to decentralized social prediction tasks, which provides new capabilities for building more personalized, user-driven, and decentralized social network systems.
To overcome this gap, we propose DeSocial framework that allows users to regain control over the forecasting process on their own behalf. In DeSocial , each user has their rights to select the most favorable prediction model from a library of graph learning networks (e.g., MLP [51], GCN [24], GAT [55], GraphSAGE [13], or SGC [64]). This selection is based on neighborhood sampling evaluation to ensure that the selected model best captures the user’s local environment. To ensure reliability, DeSocial applies a smart contract that is transparent to every nodes: a group of validators are selected to run their models independently, evaluate the link prediction query, and make the decision together by a majority vote. This approach not only increases the robustness of the predictions, but also removes single-point algorithmic control, aligning outcomes with both user intent and system integrity. Our contributions can be summarized as follows:
@@ 2.1 Graph Learning for Social Networks
Existing graph learning methods for social networks are advanced, including graph neural networks (GNN) [24, 55, 13, 64, 71, 48, 14, 26, 52, 43, 59], meta-learning frameworks [91, 6, 83, 77, 57], Transformer-based [66, 65, 82], and large language model (LLM)-based [34, 53, 4, 60, 80, 15, 85, 92], most of which are designed for centralized environments, where the full graph structure and node features are aggregated on a single server and optimized jointly. In contrast, our setting assumes a completely decentralized environment where each user node has full access to the local view of the social network structure, but raw data cannot be shared across nodes [87, 45, 37, 46]. Furthermore, the storage and computational capabilities of each node are limited, making it infeasible to deploy large-scale models such as LLM-based encoders [28, 53, 20, 49] or even invoke external LLM APIs [73, 33, 36, 74, 89, 86, 22, 73, 23, 21]. This practical limitation leads us to limit each node to implementing a lightweight graph foundation. Therefore, we selected five representative and well-studied models: MLP [51], GCN [24], GAT [55], GraphSAGE [13], and SGC [64], as candidate backbones for our framework. These backbones achieve a practical balance between representational power and computational efficiency, which is sufficient to test the effectiveness of model selection and consensus in decentralized graph learning.
@@ 2.2 Blockchain Consensus Mechanism
Blockchain consensus mechanisms enable distributed agreement without central control [39, 42, 63]. The original Bitcoin blockchain relied on Proof-of-Work (PoW) [42] consensus, where the longest chain determines the accepted state. While PoW provides an open network, it is energy intensive and only provides probabilistic finiteness. More relevant to our context are the consensus protocols Proof-of-Stake (PoS) [63, 76, 2], Proof-of-Authority [1, 2], and Byzantine Fault Tolerance (BFT) [27, 35, 54], which achieve finality by voting among selected validators. However, these methods do not achieve complete decentralization [9, 18], or a large number of validators are selected to participate in the verification, which greatly reduces the operation speed of the network [81]. These concerns of scalability, centralization, and running efficiency bring challenges to blockchain-based decentralized social network algorithm design. Therefore, a limited set of validators are selected to participate in the consensus for each verification in DeSocial , which guarantees the efficiency, robustness, and complete decentralization of the network.
@@ 2.3 Ensemble Learning and Majority Voting
Ensemble learning that combines multiple models to improve predictions is well-recognized in machine learning [7, 3, 11, 38, 50]. In graph learning, ensemble methods help mitigate overfitting to specific topologies or noisy subgraphs, especially when models are trained on different views of the graph or initialized with different parameters [8, 90, 5]. Voting is essential in blockchain consensus process because it enables nodes to collectively agree on the validity of transactions [70]. Similarly, voting serves as a mechanism for aggregating decisions from multiple nodes in decentralized graph learning, ensuring that predictions are agreed upon through collective decision-making rather than centralized authority. Common voting strategies includes soft voting [61, 79], where models make probabilistic predictions and use averages, and hard voting [10, 61, 79], where each model casts a binary vote and the final outcome is the majority. DeSocial adopts hard voting since it aligns naturally with blockchain consensus mechanisms, where validations are made through majority approval. Voting-based decisions are also more robust and less vulnerable to manipulation.
@@ 5.1 Experimental Setups
We use four real-world temporal graph datasets in the scope of Web 2.0 (UCI [44], Enron [88], and GDELT [88]) and Web 3.0 (Memo-Tx [94]). In order to evaluate the performance of DeSocial , we used the five most classic centralized models (MLP [51], GCN [24], GAT [55], GraphSAGE [13], and SGC [64]) in the field of social network learning as baselines for comparison. Detailed information for the datasets and baselines is described in Appendix  A.1 and Appendix  A.2, respectively.
@@ 5.2 Evaluation Metrics
Given the decentralized nature of our problem formulation, we employ evaluation metrics tailored to assess the effectiveness of DeSocial in decentralized environments, enlightened by the validation of the double spending problem on blockchain [32, 25]. That is, they lack access to the underlying intent of a transaction and can only judge based on structural context. This motivates our use of Acc@K, a set of relative metrics that evaluates whether the predicted positive link ranks higher than its sampled alternatives.
Given a link prediction task that each test case consists of one positive edge and $K-1$ negative edges, Acc@K indicates the probability that the model assigns the highest score to the positive edge among the $K$ candidates. We adopt $K\in\{2,3,5\}$ to evaluate the performance of decentralized graph learning. The greater $K$ is, the harder evaluation tasks will be. We follow  [88] for the randomized negative sampling method. Appendix  A.3 shows our experiment setups and implementation details.
@@ 5.3 Impacts of the Personalized Algorithms
Figure 4 reports the performance of five centralized baselines, random selection, simple rule-based selection, and our personalized algorithm selection method DeSocial PA across multiple datasets and evaluation metrics. In hybrid settings, each validator uses a distinct backbone. Random selection represents that every node selects an algorithm in $\mathcal{F}$. Simple rule-based selection represents that every node selects an algorithm in $\mathcal{F}$ based on its two features (degree and clustering) of the local structure.  Table 1 and  Table 2 shows the detailed statistics of the centralized baselines and DeSocial PA. Appendix  A.5 shows the simple rule of algorithm selection.
@@ 5.5 Efficiency Analysis
We use Ganache to capture blockchain-side latency and execution cost in decentralized social prediction processes, avoiding the complexity of deploying to a full public ETH testnet. For decentralized methods, runtime is measured from a single user’s perspective. That is, starting from a social link request and ending to the decision aggregated. Appendix  B.2 shows the details of calculating the run time of each decentralized method.
As shown in  Table 3, the amortized runtime of DeSocial is not significantly impacted by the integration of the ETH local development chain. Our experiments are conducted in a single machine simulation where both graph training and blockchain interactions are executed serially. Appendix  A.3 shows the details of the computation device. In a real ETH network deployment with multiple participating users, the system would better exploit parallelism and further improve the efficiency.